# Unsupervised full-field Bayesian inference of orthotropic hyperelasticity from a single biaxial test: a myocardial case study

### v1

@@ 6 Validation
We reproduce multimodal, homogeneous loading protocols, i.e. classical incompressible biaxial tension and triaxial shear tests, following the experimental protocols of (Dokos et al., 2002; Sommer et al., 2015). Constitutive parameters inferred from our unsupervised single-shot heterogeneous biaxial protocol are then used to assess, in a controlled fashion, the predictive capability of the proposed framework relative to established multimodal calibration pipelines for myocardial tissue (Guan et al., 2018; Peirlinck et al., 2018; Kakaletsis et al., 2021; Martonová et al., 2024). Specifically, we prescribe homogeneous deformations with deformation gradient:
For the biaxial tension cases, the fiber and normal stretches $\lambda_{\rm f}$ and $\lambda_{\rm n}$ are controlled, the sheet stretch follows from incompressibility, $\lambda_{\rm s}=1/(\lambda_{\rm f}\lambda_{\rm n})$, and all shear strains are set to zero, $\gamma_{\rm{fs}}=\gamma_{\rm{sf}}=\gamma_{\rm{nf}}=\gamma_{\rm{fn}}=\gamma_{\rm{sn}}=\gamma_{\rm{ns}}\equiv 0$. The hydrostatic pressure $p$ is determined by enforcing zero normal traction in the sheet direction, $\sigma_{\rm ss}=0$, which yields the following non-zero normal stresses:
[…]
Here $\psi_{\star}$ indicates the partial derivative of the strain energy density function with respect to the corresponding invariant $I_{\star}$, i.e. $\partial\psi/\partial I_{\star}$. In the triaxial shear tests, the principal stretches are fixed to unity, $\lambda_{\rm f}=\lambda_{\rm s}=\lambda_{\rm n}\equiv 1$, and only one shear component is activated per test; all others are set to zero, $\gamma_{\rm{fs}}=\gamma_{\rm{sf}}=\gamma_{\rm{nf}}=\gamma_{\rm{fn}}=\gamma_{\rm{sn}}=\gamma_{\rm{ns}}\equiv 0$, with a single component allowed to vary. Each test therefore excites a single shear mode. Given that $\gamma_{\star}\geq 0$, the resulting shear stresses are:
Figure 10: Noise-free validation single-shot versus multi-modal biaxial and triaxial shear tests. Comparing Cauchy stresses based on the ground-truth parameter values, shown in tri markers, and the inferred noise-free full-field single-shot biaxial test parameter values sets, shown in shaded colors representing the $95\%$ confidence intervals resulting from the inferred posterior parameter set distributions. $R^{2}$ scores for inferred versus ground-truth stress values, for each individual biaxial tensile and triaxial shear test, are shown in the top left corners.
@@
Figure 10 reports noise-free validation across three heterogeneous configurations: circ–long–het (left), circ–rad–het (middle), and rad–heli–het (right). Each column compares Cauchy stresses from ground-truth parameters (tri markers) with predictions from the single-shot, noise-free full-field inference constitutive parameter sets (shaded 95% credible intervals) under classical biaxial tension and triaxial shear protocols. In the circ–long–het configuration, the inferred biaxial stress–stretch curves match the ground-truth with $R^{2}=1.00$ for both biaxial tension cases. The nf-, sn-, and ns-shear responses are also indistinguishable from the ground-truth, whereas the cross-fiber shear components show lower agreement, with $R^{2}=0.89$ for $\sigma_{\mathrm{fs}}$ and $R^{2}=0.77$ for $\sigma_{\mathrm{sf}}$ respectively. Aggregated over all tests in this configuration, the overall coefficient of determination is $R^{2}=0.96$. These deviations are consistent with limited excitation of the $I_{4fs}$ shear-coupling invariant in the circ–long-het configuration as shown in Figure 5. In the circ–rad–het configuration, the inferred constitutive parameter sets reproduce the ground-truth stresses across biaxial tension and all shear tests with $R^{2}=1.00$ throughout. Posterior confidence intervals remain tight, indicating precise recovery of the active parameters for this orientation. In the rad–heli–het configuration, inferred parameter set curves overlay the ground-truth for all biaxial tensile and shear tests with $R^{2}=1.00$. Overall, the noise-free validation shows that single-shot full-field inference generalizes across all three heterogeneous configurations, with slight deviations confined to cross-fiber shear in the circ–long–het case.
Figure 11: High-noise validation single-shot versus multi-modal biaxial and triaxial shear tests. Comparing Cauchy stresses based on the ground-truth parameter values, shown in tri markers, and the inferred high-noise $\sigma=10^{-3}$ full-field single-shot biaxial test parameter values sets, shown in shaded colors representing the $95\%$ confidence intervals resulting from the inferred posterior parameter set distributions. $R^{2}$ scores for inferred versus ground-truth stress values, for each individual biaxial tensile and triaxial shear test, are shown in the top left corners.
@@
Figure 11 reports high-noise ($\sigma=10^{-3}$) validation across three heterogeneous configurations: circ–long–het (left), circ–rad–het (middle), and rad–heli–het (right). Each column compares Cauchy stresses from ground-truth parameters with predictions from the single-shot full-field inference under displacement noise $\sigma=10^{-3}$ (shaded 95% confidence intervals). In the circ–long–het configuration, inferred biaxial tension behavior remains perfectly aligned with the ground-truth stress behavior. Similarly, the nf-, sn-, and ns-shear responses retain correct magnitude and slope. Noise primarily affects the inferred cross-fiber shear stress $\sigma_{\mathrm{fs}}$ and $\sigma_{\mathrm{sf}}$ response, where credible intervals widen and small deviations between ground-truth and inferred behaviors appear. Aggregated over all tests in this configuration, the overall coefficient of determination is $R^{2}=0.96$. In the circ–rad–het configuration, the stresses computed from the inferred parameter sets under noisy full-field biaxial tensile tests, no longer fit the ground-truth stress curves perfectly. Mean trends are preserved across biaxial tensile and shear protocols, but panel-wise $R^{2}$ values are lower than those in the high-noise circ-long-het and rad-heli-het configurations. Nevertheless, the aggregated coefficient of determination for this configuration still remains $R^{2}=0.77$. In the rad–heli–het configuration, we see the best overall inferred stress responses across all biaxial tensile and shear stress responses, with $R^{2}$-values greater than 0.88 across the board. Here, the total aggregated coefficient of determination equals $R^{2}=0.93$. Overall, despite elevated displacement and thus deformation noise, the single-shot SVI inference approach maintains very good quantitative agreement with ground-truth stresses across all three heterogeneous configurations.
@@ 7 Discussion
The ability to infer complex orthotropic hyperelastic models from a single full-field biaxial tensile testing experiment represents an interesting avenue in biomechanical tissue characterization. Traditional workflows typically rely on multiple specialized loading configurations and assume simplified model forms to ensure parameter identifiability (Dokos et al., 2002; Sommer et al., 2015; Avazmohammadi et al., 2018). In contrast, our framework leverages specimen-intrinsic microstructural (and geometrical) heterogeneity, combined with stochastic variational inference, to robustly recover full sets of nonlinear, cross-correlated material parameters under realistic noise conditions. This approach not only reduces experimental tissue manipulation but also expands opportunities to characterize spatially heterogeneous biological tissues, where sample availability or repeat testing is limited. Moreover, it lays the groundwork for practical integration of physics-informed Bayesian inference into future experimental biomechanics and computational modeling pipelines. Additionally, we offer the possibility of taking inter-sample and intra-sample variability into account for materials where this was previously not an option.
Power of tissue testing heterogeneity in constitutive parameter identifiability. Our results show that introducing both geometric and microstructural heterogeneity in the training specimen substantially improves the identifiability of constitutive model parameters. The addition of a simple circular occlusion generates spatial strain gradients and local stress concentrations that increase the sensitivity of internal forces to variations in model parameters (Attar et al., 2022; Conde et al., 2023; Pottier et al., 2011; Pierron and Grédiac, 2021). Similarly, strategically aligning fiber orientations relative to the applied boundary conditions enhances the excitation of multiple deformation invariants (Zhang et al., 2022; Kakaletsis et al., 2021; Linka et al., 2021). These experimental design choices enabled stable recovery of all parameters associated with dominant energy modes ($I_{1}$, $I_{4ff}$), and even parameters associated with more subtle contributions ($I_{4fs}$) when measurement noise was low. Importantly, both forms of heterogeneity complemented each other: geometric features enriched the macroscopic stress states, while fiber orientation tuning ensured multiaxial activation of the microstructure. These findings suggest that future experimental protocols could deliberately incorporate both microstructural and geometrical tailoring to maximize parameter observability.
Homogeneous loading conditions heavily impact orthotropic inference power. Samples with homogeneously distributed fiber orientations and without geometric features exhibited markedly worse parameter convergence (Pottier et al., 2011). In these homogeneous specimens, the dominant loading paths aligned primarily with the $I_{1}$ and $I_{4ff}$ invariants, while parameters associated with fiber-sheet coupling ($I_{4fs}$) received little excitation, as previously postulated by (Holzapfel and Ogden, 2009; Guan et al., 2018). As a result, posterior distributions for associated parameters such as $a_{fs}$ and $b_{fs}$ remained close to their prior, showing truncation effects and large uncertainty. Interestingly, under high measurement noise, homogeneous configurations experienced smaller decrease in performance in validation compared to heterogeneous specimens. This reflects a robustness trade-off: simpler configurations concentrate strain energy into a few modes, reducing sensitivity to noise at the cost of losing full identifiability. These findings highlight that while homogeneity can be beneficial under extreme noise, achieving accurate full-model inference requires careful experimental enrichment.
Measurement noise, as expected, degrades inference performance. At a displacement noise level of $1\times 10^{-4}$ mm, the method still achieved low relative errors and tightly concentrated posterior distributions for most constitutive parameters. The shear related parameter $a_{fs}$ and $b_{fs}$, however, show larger deviation from the ground-truth. At $1\times 10^{-3}$ mm noise, significant broadening and bias in the posterior appeared, particularly for parameters linked to minor strain energy contributions. Namely, the fourth coupling invariant $I_{4fs}$ showed strong vulnerability to noise, with associated parameters $a_{fs}$ and $b_{fs}$ tending toward zero or staying close to the prior resulting in a truncated posterior distribution, which is in agreement with the loss in accuracy for anisotropic features in (Rahmani et al., 2013). This effectively removed their contribution to the strain energy function in the inferred models. Despite these deviations, the validation strain energy and stresses remained close to the ground-truth. This suggests that some parameter errors can be tolerated without substantial loss of predictive accuracy for macroscopic mechanical behavior. Improving noise robustness through enhanced experimental design, probabilistic displacement modeling, or variational reweighting strategies offer important future directions. (Marek et al., 2019; Jafari et al., 2025; Peyraut and Genet, 2025)
Single-shot inferred material models validate well to multimodal testing protocols - even under high noise. We show that we can reproduce the behavior from classical homogeneous multimodal biaxial tensile and triaxial shear testing. Even at high noise levels, we see better overall agreement with the ground-truth stresses from a single heterogeneous full-field biaxial tensile test than we generally see for separate biaxial tensile-only, or triaxial shear-only training. Interestingly, the circ-rad-het configuration seems to be most sensitive to displacement noise when considering pure biaxial tensile and triaxial shear test loading, despite this configuration’s richer heterogeneous deformation profile shown in Figure 5. Both the circ-long-het and rad-heli-het samples show good extrapolative behavior, even under high noise levels when the shearing component in the strain energy vanishes, forming interesting slicing configurations for future experimental studies of myocardial tissue behavior.

### v2

@@ 5.1 Uni-modal deformation mode validation
We reproduce homogeneous, uni-modal loading protocols, i.e. classical incompressible biaxial tension and triaxial shear tests, following the experimental protocols of [10, 11]. We consider incompressible biaxial tension, where $\lambda_{\rm f}$ and $\lambda_{\rm n}$ are prescribed and $\lambda_{\rm s}$ follows from incompressibility, and triaxial shear tests, where the principal stretches are fixed to unity and a single shear component is activated per test. The resulting stress expressions are summarized in D. Constitutive parameters inferred from our unsupervised single-shot heterogeneous biaxial protocol are then used to assess, in a controlled fashion, the predictive capability of the proposed framework relative to established multi-modal calibration pipelines for myocardial tissue [12, 15, 16, 26].
Figure 10: Noise-free single-shot inference validation under homogeneous uni-modal loading. Cauchy stresses predicted by the ground-truth HO parameter set (tri markers) and by the inferred noise-free single-shot full-field posterior (shaded $95\%$ credible intervals) for three heterogeneous configurations: circ–long–het (left), circ–rad–het (middle), and rad–heli–het (right). Rows correspond to classical incompressible biaxial tension and triaxial shear modes. Panel-wise $R^{2}$ values quantify agreement between posterior mean predictions and ground-truth stresses. Noise-free inference generalizes perfectly across all modes for circ-rad-het and rad-heli-het ($R^{2}\approx 1$ throughout), while circ-long-het shows the largest deviations in cross-fiber shear ($\sigma_{\mathrm{fs}}$, $\sigma_{\mathrm{sf}}$).
@@
Figure 10 reports noise-free validation across the three geometrically heterogeneous configurations: circ–long–het (left), circ–rad–het (middle), and rad–heli–het (right). Each column compares Cauchy stresses from the ground-truth parameter set (tri markers) with predictions based on the noise-free single-shot full-field inference parameter sets (shaded $95\%$ credible intervals) under classical biaxial tension and triaxial shear tests. Leveraging parameters inferred from the noise-free circ-long–het slab, the inferred biaxial stress-stretch curves match the ground-truth mechanical response of the tissue with $R^{2}=1.00$ for both biaxial tension cases. The inferred nf-, sn-, and ns-shear responses are also indistinguishable from the ground-truth response, whereas the inferred cross-fiber shear behavior shows lower agreement with respect to the ground-truth model, with $R^{2}=0.89$ and $R^{2}=0.77$ for $\sigma_{\mathrm{fs}}$ and $\sigma_{\mathrm{sf}}$ respectively. Aggregated over all tests in this configuration, the overall coefficient of determination amounts to $R^{2}=0.96$. These deviations are consistent with the limited deformational excitation of the $I_{4fs}$ shear-coupling invariant in the circ-long-het specimen (Figure 5) and with the weaker identifiability of $(a_{fs},b_{fs})$ observed in the previous section. Leveraging the circ-rad-het slab, the inferred constitutive parameter sets perfectly reproduce the ground-truth stresses across all biaxial tension and shear tests with $R^{2}=1.00$ in every panel. Posterior confidence intervals remain tight, indicating precise recovery of the underlying constitutive parameters for this orientation, including the shear-coupling response. Focusing on the rad-heli-het specimen, the inferred stress curves again overlay the ground-truth for all biaxial tensile and shear tests with $R^{2}=1.00$. Overall, the noise-free validation shows that single-shot full-field inference generalizes reliably across all three heterogeneous configurations, with slight deviations confined to cross-fiber shear in the circ-long-het case, which is in agreement with the parameter-level identifiability analysis in Figure 7.
Figure 11: High-noise single-shot inference validation under homogeneous uni-modal loading. Cauchy stresses predicted by the ground-truth HO parameter set (tri markers) and by the inferred single-shot full-field posterior under displacement noise $\sigma_{u}=10^{-3}\,\mathrm{mm}$ (shaded $95\%$ credible intervals) for circ–long–het (left), circ–rad–het (middle), and rad–heli–het (right). Rows correspond to the same biaxial tension and triaxial shear protocols as in Figure 10. Panel-wise $R^{2}$ values quantify agreement between posterior mean predictions and ground-truth stresses. High noise primarily degrades cross-fiber shear predictivity and widens credible intervals, most notably for circ-rad-het, while biaxial tension and non-cross-fiber shear modes remain well captured, particularly for circ-long-het and rad-heli-het.
@@
Figure 11 reports high-noise ($\sigma_{u}=10^{-3}$) validation for the same three heterogeneous single-shot slab specimens. Inferring from displacement and reaction force data in the circ–long–het specimen, the inferred biaxial tension behavior remains well aligned with the ground-truth, with panel-wise $R^{2}$ values close to one. Similarly, the nf-, sn-, and ns-shear responses retain correct magnitude and slope. Noise primarily affects the inferred cross-fiber shear stresses $\sigma_{\mathrm{fs}}$ and $\sigma_{\mathrm{sf}}$, where credible intervals widen and small systematic deviations between ground-truth and inferred curves appear. Aggregated over all tests in this configuration, the overall coefficient of determination remains high at $R^{2}=0.96$. Leveraging the circ–rad–het slab, stresses computed from the high-noise inferred parameter sets no longer match the uni-modal ground-truth responses perfectly. Mean trends are preserved across biaxial and shear protocols, but panel-wise $R^{2}$ values are lower than in the circ-long-het and rad-heli-het configurations, particularly in cross-fiber and sheet-direction shear. Nevertheless, the aggregated coefficient of determination still reaches $R^{2}=0.77$. Lastly, we note the best uni-modal predictive nature of the inferred parameter set from the high-noise rad-heli-het specimen, with with all panel-wise $R^{2}$-values greater than $0.88$ and an aggregated coefficient of determination of $R^{2}=0.93$. Overall, despite elevated displacement noise and the associated degradation in certain parameter estimates reported in Section 4.2, the single-shot stochastic variational inference framework maintains very good quantitative agreement with uni-modal ground-truth stresses across all three heterogeneous configurations, especially in the biaxial tension and non-cross-fiber triaxial shear modes.
5.2 Mixed-modal deformation validation
To probe generalization under mixed deformation modes, we sample deformation states in invariant spaces between the homogeneous loading paths. We draw $(I_{1},I_{4ff},I_{4nn},I_{4fs})$ uniformly within the minimum and maximum values induced by the homogeneous protocols in the uni-modal deformation, see also (31), and retain only physiologically admissible states. Specifically, we accept samples that satisfy near-incompressibility $0.9999\leq J\leq 1.0001$ and whose maximum principal Cauchy stress does not exceed $20.3~\mathrm{kPa}$, consistent with reported peak stresses for left-ventricular biaxial tests on human myocardial tissue [11]. The invariant-to-kinematics reconstruction and stress evaluation used for filtering are detailed in E. We repeat this procedure until 5,0005{,}000 admissible mixed-modal states are obtained.
Figure 12: Mixed-modal validation set construction in invariant space. Invariant trajectories induced by uni-modal biaxial and single-shear protocols (colored curves) compared with randomly sampled invariant combinations that satisfy (quasi-)incompressibility and a physiological bound on maximum principal stress detailed in E. The uni-modal protocols excite low-dimensional paths, whereas the accepted multi-modal deformation modes populate the space between these uni-modal paths, yielding mixed-modal deformation states for further and richer out-of-training validation.
Figure 12 illustrates the resulting invariant coverage. The thin sweeping colored arcs represent the uni-modal loading conditions from the previous section, highlighting that these tests excite the invariants along low-dimensional paths. In contrast, our randomly sampled deformation gradients populate the space between these arcs and thus correspond to more complex mixed-modal testing protocols. The accepted samples lie between $3$ and $3.25$ in $I_{1}$, between $0.57$ and $1.32$ for $I_{4ff}$ and $I_{4nn}$, and between 0.5-0.5 and $0.5$ for $I_{4fs}$, and they are spread out approximately uniformly within this constrained region. This construction supports the view that the resulting invariant set spans a physiologically relevant range of deformation states that is well suited for out-of-training validation of the inferred versus ground-truth constitutive parameter set.
Figure 13: Strain-energy generalization under mixed-modal deformations. Correlation between strain energy predicted by inferred parameter posteriors, $\Psi$, and the ground-truth strain energy, $\Psi^{\mathrm{GT}}$, over 5,0005{,}000 accepted mixed-modal deformation states sampled in invariant space (Figure 12). Results are shown for geometrically heterogeneous training specimens under both low and high displacement measurement noise, i.e. $\sigma_{u}=10^{-4}$ and $10^{-3}\,\mathrm{mm}$ respectively. Hexbin density emphasizes the dominant response region, and the black diagonal indicates perfect agreement. Despite noise-driven degradation of select parameter groups, predicted energies remain strongly correlated with the ground-truth across mixed-modal states (all $R^{2}\geq 0.9$).
More specifically, we compute strain energies using both the ground-truth parameter set and the inferred parameter sets obtained from both low and high displacement measurement noise-affected heterogeneous training experiments. The correlation between the strain energy predicted by the inferred parameter sets $\Psi$ and the ground truth $\Psi^{\mathrm{GT}}$ is shown as a hexbin plot in Figure 13. The black diagonal centerline indicates optimal behavior, i.e. the inferred mechanical response matches the expected ground-truth response perfectly. In line with our prior results for the weakest identifiability of the noise-affected circ-long–het specimen in the parameter inference and uni-modal stress inference analysis, this specimen also yields the lowest strain-energy agreement across the multi-modal deformation modes with $R^{2}=0.91$ in the low-noise setting and $R^{2}=0.931$ in the high-noise setting. Across the three heterogeneous slabs tested, we observe larger deviations away from the diagonal, indicating that a subset of randomly sampled invariant combinations, together with corresponding posterior samples, lead to strain energies that differ more strongly from the ground-truth. The circ-rad-het specimen shows a very high coefficient of determination in the low-noise case ($R^{2}=0.989$), and its performance slightly degrades to $R^{2}=0.915$ in the high-noise case. Despite this drop, circ-rad-het exhibits the tightest clustering of hexbin intensities around the diagonal, suggesting that most mixed-modal states are still well captured, with a relatively small number of outliers contributing to the $R^{2}$ reduction. Lastly, the most consistent performing sample is the rad-heli-het with an $R^{2}$ score of $0.962$ in the low noise setting and $0.968$ in the high noise setting.
Overall, the mixed-modal strain-energy validation indicates that, even though high displacement noise can substantially perturb individual parameter estimates—particularly the shear-coupling and some directional anisotropic terms—the inferred models retain strong predictive capability at the level of strain energy across a broad set of physiologically admissible deformation states (all $R^{2}\geq 0.9$). This finding complements the stress-based validation and suggests that the proposed single-shot full-field identification framework can generalize reliably beyond the specific training protocol, provided that geometric and microstructural heterogeneity are sufficiently rich and measurement noise remains within a moderate range.
@@ 6 Discussion
The ability to infer complex orthotropic hyperelastic models from a single full-field biaxial tensile testing experiment represents an interesting avenue in biomechanical tissue characterization. Conventional constitutive identification workflows for complex orthotropic tissues, suchs as the heart muscle, typically rely on assumed microstructural homogeneity across the tissue specimen and leverage multiple specialized uni-modal loading protocols to identify constitutive parameter sets [10, 11, 74]. This reflects a broader challenge for orthotropic HO myocardium laws: parameter uniqueness is generally difficult to guarantee under limited excitation, motivating multi-modal protocols, sensitivity-informed reparameterizations, or model reduction [15, 75, 76, 77] This work demonstrates that a single heterogeneous biaxial experiment, when combined with full-field kinematics and sparse reaction-force measurements, can support uncertainty-aware inference of a nonlinear orthotropic Holzapfel–Ogden (HO) parameter set using stochastic variational inference. Beyond the identifiability perspective, this single-shot strategy reduces experimental tissue manipulation and is particularly attractive for heterogeneous biological tissues, where sample availability or repeat testing is limited. Across the considered slab configurations, two consistent findings emerge: (i) identifiability is primarily governed by the degree to which the experiment excites the relevant invariant subspace, and (ii) even when individual HO parameters are not recovered accurately under varying levels of displacement measurement noise, the inferred models can remain highly predictive for stresses and strain energy over both uni-modal and mixed-modal deformation states. These observations clarify when single-shot full-field inference can be expected to recover the full parameter set, and when it should instead be interpreted as delivering a reliable posterior predictive constitutive response with quantified uncertainty.
Heterogeneity as an experimental design principle for orthotropic constitutive parameter inference. Dominant gains in constitutive parameter identifiability arise from deliberately enriching the deformation space within a single test. Microstructural heterogeneity, introduced through slab orientation inducing specific within-specimen heterogeneity through transmural fiber rotation, and geometric heterogeneity, introduced through a central occlusion, both broaden the activated distributions of $I_{1}$, $I_{4ff}$, $I_{4nn}$, and $I_{4fs}$ (Figures 4-5). These microstructural or geometric experimental protocol choices increase the sensitivity of the equilibrium residuals to the full parameter vector $\bm{\theta}$. In the noise-free setting, this explains the contrast between homogeneous and heterogeneous specimens in Figures 6-7. Most isotropic and primary anisotropic parameters are inferred robustly across configurations, while the shear-coupling parameters $(a_{fs},b_{fs})$ become practically identifiable only when the single slab biaxial tension experiment produces appreciable $I_{4fs}$ excitation. The circ-rad-het and rad-heli-het slab configurations provide the most consistent recovery of these terms, reflecting the combined effect of geometry-induced strain gradients and orientation-driven multiaxial activation of the microstructure. In this sense, our results translate the multi-modal calibration message from prior myocardium studies into a single-shot setting: rather than combining separate stretch and shear protocols, we aim to excite the same (otherwise weakly constrained) coupling modes within one heterogeneous experiment [10, 11, 15]. These results further suggest that future protocols could deliberately combine geometrical and microstructural tailoring to maximize parameter observability within a single specimen.
Experimental displacement measurement noise selectively impacts constitutive parameter identifiability. Displacement noise degrades constitutive parameter identifiability selectively rather than uniformly. For low noise ($\sigma_{u}=10^{-4}$ mm), Figures 8-9 show stable convergence and comparatively narrow posteriors for the isotropic and primary anisotropic parameters. In contrast, $(a_{fs},b_{fs})$ shows the largest posterior broadening and, in some cases, pronounced skewness or truncation of the inferred posterior parameter distribution. For high noise ($\sigma_{u}=10^{-3}$ mm), the shear-coupling terms lose practical identifiability across specimens, and additional degradation can appear in directional anisotropic parameters, most clearly in circ-rad-het. These results are consistent with global sensitivity and inverse uncertainty quantification studies of the underlying Holzapfel-Ogden constitutive model [43, 77, 78], which show that parameter sensitivity and practical identifiability depend strongly on the deformation modes and invariant support represented in the available data. Two implications follow. Full-field single-shot inference strongly benefits from experimental designs that amplify invariants associated with weaker energy modes. In addition, posterior-shape diagnostics such as truncation and skewness provide useful indicators of noise-driven practical non-identifiability and should be interpreted alongside point estimates.
Identifiability versus predictive capability: how important is it to identify all the underlying constitutive parameters. An important result is that imperfect recovery of ground-truth HO parameters does not necessarily imply poor constitutive prediction, as also noted in prior HO calibration and in vivo estimation studies [75]. Figures 10-11 show that posterior predictive stresses under classical homogeneous uni-modal protocols remain accurate in aggregate, even at high noise. Figure 13 shows strong strain-energy correlation across 5,000 physiologically admissible mixed-modal deformation states, with all $R^{2}\geq 0.9$. When viewed together with the parameter-level degradation in Figures 8-9, these results indicate that distinct parameter combinations can yield near-indistinguishable macroscopic responses over the rich explored invariant deformation space [15, 77] In such cases, the posterior predictive response, together with its credible interval becomes more relevant for downstream simulation than any single best-fit parameter vector and is consistent with inference perspectives that emphasize parameter ranges rather than unique point estimates. This perspective is particularly pertinent for soft tissues, where measurement noise, limited loading diversity, and intrinsic variability constrain practical identifiability.
Some heterogeneous configurations remain more predictive under noise. Our validation results also reveal slab-specific trends. The circ-rad-het configuration can be more sensitive to displacement noise in the uni-modal and mixed-modal validations, despite exhibiting a rich heterogeneous deformation profile in Figure 5. This observation indicates that heterogeneity alone is not sufficient. Identifiability also depends on how displacement measurement noise perturbs the invariants that control weaker energy modes and how these perturbations propagate through the equilibrium residuals, which is consistent with the broader observation that the subset of HO terms supported by data depends on the deformation modes represented in the calibration dataset [15]. In contrast, rad-heli-het exhibits the most consistent predictive performance across stress and energy validations in Figures 10-13, making it a promising slicing orientation when a single-shot protocol is prioritized and robust in-plane excitation of multiple microstructural constituent contributions is feasible.
Outlook. Some limitations motivate future work. The results are obtained on synthetic data with known boundary conditions, known microstructural fields, and an assumed HO model form. Future experimental work will need to address boundary-condition uncertainty, microstructure dispersion error or misregistration, and model discrepancy [79, 80]. More broadly, automated model discovery frameworks provide an interesting avenue to deal with both practical non-identifiability and model-form discrepancy in settings with complex loading patterns and varying amounts of measurement noise [16, 30, 32, 33, 53, 81] Complementary directions include explicit probabilistic models for displacement measurement error, experimental designs that target weakly activated invariants, and alternative variational weighting strategies to improve robustness under noisy full-field data [29, 73, 82]. Variational families that capture posterior dependencies more accurately, combined with posterior predictive checks tailored to equilibrium residuals, would further strengthen uncertainty quantification. Finally, the present study focused on spatially homogeneous material parameters with spatially varying structure fields. Extending the framework to spatially varying parameter fields with appropriate regularization could be valuable for truly heterogeneous tissues, while raising additional identifiability and experimental design challenges [83, 84].
Overall, our results support single-shot, uncertainty-aware identification of nonlinear orthotropic constitutive behavior from a single heterogeneous biaxial test. Our main experimental design message is that the relevant invariant subspace must be sufficiently activated, with particular emphasis on the weak shear-coupling mode. At the same time, the validation results show that posterior predictive accuracy can remain high even when individual parameters are noise-limited.